A rope pulls upwards a bucket of water of mass 15kg from a shallow well. The bucket accelerates at 2.2m/s2. what is the tension on the rope


1
Expert's answer
2021-02-22T16:03:03-0500
"T=ma+mg=m(a+g)\\T=15(2.2+9.8)=180\ N"